OpenStudy (anonymous):

in a survey with 50 members, 19 take math, 35 take english, and 9 take both. how many take neither?

OpenStudy (anonymous):

i still need this problem solved immediately please can someone help me out!!!!

OpenStudy (solomonzelman):



OpenStudy (solomonzelman):

just math. 19-9=10 just English 35-9=26 both =9 so all of the takers, 10+26+9=45 There are 50 people, and 45 of them are taking surveys. 50-45=5 

OpenStudy (solomonzelman):

@bboone69, see?